# Supernova Electron-Neutrino Interactions with Xenon in the nEXO Detector

### v1

@@ III.4 Comparison of inclusive cross section predictions
As a test of MARLEY’s predictions with the supplied nuclear data, cross sections for $\nu_{e}$ interactions with 134Xe and 136Xe from MARLEY are compared to the calculations in Ref. [22], shown in Fig. 1. In that reference, calculations of CC cross sections for stable isotopes of xenon are provided, including contributions from forbidden transitions. As can be seen in the figure, MARLEY agrees fairly well with those predictions at lower incident neutrino energies, but its predictions are smaller at higher energies. This could be in part due to the omission of forbidden transitions in the approximation used by MARLEY.
@@
The electron-neutrino flux-averaged cross sections are computed using three commonly used neutrino flux models: the GKVM model [50], the Livermore model [51], and a “pinched-thermal” spectrum with parameters $(\alpha,\varepsilon,\langle E_{\nu}\rangle)$ of (2.5, 5×$\times 10^{52}$ ergs, 9.5 MeV) (from Ref. [14], see Sec. V.1 for a discussion of these parameters). As the majority of the incident supernova electron-neutrino flux is expected to occur at lower energies, the flux-averaged cross sections from MARLEY agree well with the predictions from Ref. [22], as can be observed in Tab. 1. While flux-averaged cross sections vary significantly with different neutrino flux models, the normalization for the neutrino flux is different as well, so the expected interaction rates with different models are closer together (see Fig. 2).
@@
Figure 1: Comparison of MARLEY’s prediction for the inclusive electron-neutrino CC cross section (using GT strengths from Refs. [34, 35]) to the calculations in Ref. [22]. Both the matrix elements used with MARLEY and those from Ref. [22] employ a quasiparticle random-phase approximation (QRPA) approach, and have the same quenching applied for comparison, corresponding to $g_{A}=0.7$.
136Xe Cross Sections (×$\times 10^{-40}\mbox{cm}^{2}$)
Flux Model Pirinen, et al. [22] MARLEY
GKVM [50] 3.15 3.10
Livermore [51] 0.89 0.83
Pinched-thermal 0.43 0.38
134Xe (×$\times 10^{-40}\mbox{cm}^{2}$)
Flux Model Pirinen, et al. [22] MARLEY
GKVM [50] 2.49 2.68
Livermore [51] 0.63 0.67
Pinched-thermal 0.28 0.28
@@
Table 1: Comparison of the predictions for inclusive cross sections from MARLEY, using the matrix elements from Refs. [34, 35], to those from Ref. [22]. For the calculations from Ref. [22], a spline interpolation is used to evaluate the cross section within the specified flux model from the interaction threshold up to 80 MeV, whereas all MARLEY cross sections are evaluated up to 100 MeV. The pinched-thermal cross section corresponds to $\alpha=2.5$, $\varepsilon=5\times 10^{52}$ ergs, and $\langle E_{\nu}\rangle=9.5$ MeV, from Ref. [14], as described in Sec. V.1.
@@
Using the inclusive cross section predictions from MARLEY and an assumed xenon mass of 4,811 kg (from Ref. [20]) enriched to 90% in 136Xe, the predicted number of electron-neutrino interactions within nEXO is plotted as a function of distance in Fig. 2, including contributions from both $\nu_{e}$-Xe interactions as well as elastic electron-neutrino scattering. Also shown is the cumulative number of red supergiants (RSGs), candidate supernova stars, within our galaxy as a function of distance, as taken from Ref. [52]. We calculate there will nominally be one or more electron-neutrino interactions in nEXO for CCSN out to $\sim$5 to 8 kpc, depending on the supernova neutrino flux model and parameters of the CCSN. Within that distance of earth there are several hundred candidate RSGs that may produce neutrino-xenon interactions in nEXO. A 300-tonne detector of similar enrichment would be able to observe events out to significantly larger distances, with $\sim$15 to 41 events predicted at a distance of 10 kpc depending on which supernova electron-neutrino flux model is used.
@@
Figure 2: The number of supernova electron-neutrino interactions ($\nu_{e}-{}^{134}$eCC, $\nu_{e}-{}^{136}$eCC and $\nu_{e}-e^{-}$) occurring in nEXO and a 300-tonne detector of the same enrichment as a function of distance. No detector efficiencies are taken into account. The red line shows the cumulative number of RSGs from Ref. [52]. The supernova flux models correspond to those from Refs. [50, 51], and a pinched-thermal spectrum as described in Sec. V.1.
@@ III.5 Exclusive cross section predictions
When a neutrino interacts with a nucleus, various particles (predominantly neutrons and gamma rays) can be emitted as a result of the de-excitation of the product nucleus. MARLEY generates predictions not only for inclusive cross sections, but also for the various exclusive cross sections and particle distributions resulting from $\nu_{e}$-Xe interactions. Exclusive cross sections leading to bound cesium states (no nucleon emission), along with those resulting in neutron emission, are shown in Tab. 2. The dominant exclusive de-excitation channel is expected to be neutron emission, although recent measurements of CC neutrino interactions on 127I and Pb with $\sim$30 MeV neutrinos have suggested this channel may be suppressed for heavy nuclei [53, 54].
@@
Channel GKVM [50] Livermore [51] Pinched-thermal
@@
Channel GKVM [50] Livermore [51] Pinched-thermal
@@
Channel GKVM [50] Livermore [51] Pinched-thermal
@@
Table 2: Calculated exclusive electron-neutrino cross sections on xenon from MARLEY, using the same assumptions as described in the caption of Tab. 1. For neutrino-electron elastic scattering, this cross section is given per xenon atom.
@@ III.6 Predicted particles and visible energy
Using MARLEY, the fraction of incident neutrino energy transferred to various output channels is visualized in Fig. 3. This plot incorporates $\nu_{e}-{}^{134}$eCC, $\nu_{e}-{}^{136}$eCC, and $\nu_{e}-e^{-}$ interactions, with appropriate cross section weighting and isotopic abundances (90% 136Xe, 10% 134Xe). Individual distributions for these three processes can be found in the appendix.
@@
In the interactions modeled, some energy will be lost to non-visible channels (thresholds for CC interactions, neutron-binding energy, and the scattered neutrino from elastic neutrino-electron scattering interactions). Additionally, emitted neutrons and the recoiling xenon nucleus also produce scintillation light and ionization charge, although these signals are quenched so their contribution is small [55]. To approximate the energy that would be visible in a detector from electron-neutrino interactions (neglecting detector thresholds, geometric size, and efficiencies), the distribution of visible (scintillation) energy is plotted along with the incident neutrino energy in Fig. 4, using an incident electron-neutrino spectrum characterized by the GKVM supernova flux model [50]. Here, visible energy ($E_{vis}$) is defined as the energy of resulting gamma rays and electrons,
@@
Figure 4: Comparison of the energy of interacting electron neutrinos and the corresponding visible energy spectrum using the GKVM flux model [50].
@@ IV Simulation and event reconstruction
Electron-neutrino interaction events were simulated using Geant4 v10.7.2 [56, 57, 58] along with a modified version of the Noble Element Simulation Technique (NEST) 2.0.1 [59, 60]. For more details on the simulation framework, refer to Ref. [20].
@@
Second, the use of charge as an energy estimator would require a stricter fiducial-volume cut to select only charge events fully-contained in the TPC drift region. As multiple MeV-scale gammas can be emitted by a CC event as part of the cesium de-excitation process, CC events are spatially larger than the events typically of interest to nEXO. Using light as the energy estimator allows for the inclusion of events that extend into the xenon space outside the drift region (referred to as the xenon skin).
While we use light as our energy estimator, we reconstruct the position and magnitude of charge depositions to use a simulated map of the geometric dependence of light collection efficiency across the detector (see Ref. [20]) to correct for the position-dependence of light collection. The impact of both grouping a light signal with an incorrect ionization signal or using partially-contained charge depositions to apply this correction is a small increase in the energy resolution of our reconstructed light signals using this algorithm. In the future, an improved charge timing and energy reconstruction can be developed for nEXO supernova events, though for this study we choose to start with the above simplifications.
@@ IV.1 Photon reconstruction
For scintillation light reconstruction, the number of collected optical photons (OPs) was calculated based on the true interaction position, the number of produced photons, a map of the simulated geometric dependence of the photon collection efficiency, the photon detection efficiency of the SiPMs, and the correlated avalanche probability (based on values measured for prototype nEXO SiPMs [20, 61]). As in Ref. [20], dark counts and electronic noise were neglected due to their small expected contribution [62, 63]. The number of OPs was calculated in a moving 300-ns window (approximately ten times the mean decay time of the long component of xenon scintillation [64]), and if the sum in that window exceeded a threshold (in terms of collected OPs, corresponding to $\sim$100efficiency for 500 keV events within the drift region), the integrated number of OPs within that window was recorded.
SiPM saturation was not included in the reconstruction algorithm, but is not expected to be a significant concern given the large spatial distribution of the energy depositions in CC events and isotropic scintillation light production. For depositions close to a SiPM that do cause saturation, it may still be possible to use non-saturated quantities of the digitized pulse such as the tail-integral [65] to reconstruct the energy of these events, although this would be needed to be studied in more detail with hardware to quantify the uncertainties it introduces.
@@ IV.2 Charge reconstruction
The number of simulated drifted ionization electrons collected by charge tiles is calculated, including effects of electron lifetime attenuation as well as longitudinal and transverse dispersion. Similarly to the light reconstruction algorithm, a rolling window of 120 microseconds is used to sum the number of detected electrons in each individual charge channel; if these exceed the threshold of 500 electrons selected for this study, the integrated charge in that charge channel is recorded.
Noise is added to the integrated charge on each channel based on a distribution derived using an optimum filter [66, 67] for reconstruction based on an internal nEXO collaboration study. This was found to be in good agreement with the data, and produces an energy resolution near the $0\nu\beta\beta$ $Q$-value consistent with that of Ref. [20]. A small number of events produce charge on channels that exceed the 120,000 electrons used in that study, so a $1/A$ dependency was fit to the channel-integrated charge noise for lower energy charge deposits and used to extrapolate noise for higher single-channel charge deposits, where $A$ is the integral of the pulse.
The potential of having multiple clusters of charge deposition (e.g. prompt gamma ray and electron depositions plus delayed neutron captures) separated in both space and time drives the need for a clustering algorithm to group charge depositions correlated with a photon signal. Rather than using the mean location of the x- and y- strips on a tile to set a location for the clustering algorithm, the charge collected by a tile is assigned to intersecting point of channels $x_{i}$ and $y_{j}$ on that tile, $Q_{x_{i},y_{j}}$,
@@
where $Q_{x_{i}}$ is the charge on channel $x_{i}$, $Q_{y_{j}}$ is the charge on channel $y_{j}$, and $Q_{x_{tile}}$ and $Q_{y_{tile}}$ are the sum of the charge on the x- and y- channels of the tile respectively. Then, the DBSCAN algorithm [68, 69] implementation of scikit-learn [70] is used to group individual charge points to form charge clusters. The DBSCAN algorithm has two free parameters, eps, which corresponds to the maximum distance between two points in the cluster for them to be considered within the same neighborhood, and min_samples, which corresponds to the minimum number of samples in a neighborhood to consider a point to be a core point. min_samples was left at its default value of 5, and eps was set to a value of 1,000 mm for this analysis. This is a conservative value chosen empirically by comparing the number of reconstructed charge clusters to the number of reconstructed light clusters. The purpose of the clustering algorithm is not to identify individual particles produced by electron-neutrino interactions, but instead to group interactions as part of either the prompt signal or the delayed neutron capture. The optimized large value likely originates from the broad spatial distribution of the emitted gammas rays and electron from these events, as well as scatters of the emitted neutron that can occur far from the electron-neutrino interaction location.
Charge-channel saturation was not incorporated into the reconstruction algorithm. Assuming a twelve-bit digitizer is used for the charge read-out, a baseline located at one-third of the digitizer range, and a gain setting corresponding to $\sim$43 electrons-per-ADC count, it was determined that $\sim$0.04of events have at least one charge channel that would be saturated. The CC events simulated are high in energy, but typically have broad spatial distributions, owing to the production of a number of de-excitation gammas, so the saturation of an individual charge channel occurs infrequently. Additionally, as with light signals it may be possible to use unsaturated portions of charge signals when saturation is present. These assumptions are motivated by simulations of the optimal settings for the CRYO ASIC charge digitization chip planned for use in nEXO [71, 72].

### v2

@@ III.4 Comparison of inclusive cross section predictions
As a test of MARLEY’s predictions with the supplied nuclear data, cross sections for $\nu_{e}$ interactions with 134Xe and 136Xe from MARLEY are compared to the calculations in Ref. [21], shown in Fig. 1. In that reference, calculations of CC cross sections for stable isotopes of xenon are provided, including contributions from forbidden transitions. As can be seen in the figure, MARLEY agrees fairly well with those predictions at lower incident neutrino energies, but its predictions are smaller at higher energies. This could be, in part, due to the omission of forbidden transitions in the approximation used by MARLEY. In Appendix A we calculate the inclusive cross section for 132Xe as a comparison.
@@
The $\nu_{e}$-Xe flux-averaged cross sections are computed using three commonly used neutrino flux models: the Gava-Kneller-Volpe-McLaughlin (GKVM) model [50] and the Livermore model [51], both from Ref. [52], and a “pinched-thermal” spectrum with parameters $(\alpha,\varepsilon,\langle E_{\nu}\rangle)$ of (2.5, 5×$\times 10^{52}$ erg, 9.5 MeV), from Ref. [14]. The flux-averaged cross sections from MARLEY agree well with the predictions from Ref. [21], as can be observed in Table 1. While flux-averaged cross sections vary significantly with different neutrino flux models, the normalization for the electron-neutrino flux and average electron-neutrino energy differ as well, so the expected interaction rates with different models are closer together (see Fig. 2).
@@
Figure 1: Comparison of MARLEY’s prediction for the inclusive electron-neutrino CC cross section (using GT strengths from Refs. [30, 31]) to the calculations in Ref. [21]. Both the matrix elements used with MARLEY and those from Ref. [21] employ a quasiparticle random-phase approximation approach and have the same quenching applied for comparison, corresponding to $g_{A}=0.7$.
136Xe Cross Section (×$\times 10^{-40}\mbox{cm}^{2}$)
GKVM [50] Livermore [51] Pinched thermal
Ref. [21] 3.15 0.89 0.43
MARLEY 3.10 0.83 0.38
134Xe Cross Section (×$\times 10^{-40}\mbox{cm}^{2}$)
Ref. [21] 2.49 0.63 0.28
MARLEY 2.68 0.67 0.28
CCSN Neutrino Model Comparison
$\nu_{e}$ (×$(\times 10^{57})$ 1.16 3.06 3.29
$\langle E_{\nu_{e}}\rangle$ (MeV) 16.5 11.3 9.5
@@
Table 1: Comparison of the predictions for inclusive cross sections from MARLEY, using the matrix elements from Refs. [30, 31], to those from Ref. [21]. For the calculations from Ref. [21], a spline interpolation is used to evaluate the cross section within the specified flux model from the interaction threshold up to 80 MeV, whereas all MARLEY cross sections are evaluated up to 100 MeV. The pinched-thermal cross section corresponds to $\alpha=2.5$, $\varepsilon=5\times 10^{52}$ erg, and $\langle E_{\nu}\rangle=9.5$ MeV, from Ref. [14], as described in Sec. V.1. Also shown is the integrated number of $\nu_{e}$s produced over the 10-sec burst and average $\nu_{e}$ energy in the various models.
@@
Using the inclusive cross section predictions from MARLEY and an assumed xenon mass of 4,811 kg (from Ref. [19]) enriched to 90% in 136Xe, the predicted number of electron-neutrino interactions within nEXO is plotted as a function of distance in Fig. 2, including contributions from both $\nu_{e}$-Xe and $\nu_{e}-e^{-}$. Also shown is the cumulative number of candidate red supergiants (RSGs), progenitors of supernovae, within our Galaxy as a function of distance from Earth, from Ref. [53]. As noted in Ref. [53], the identified set of candidate RSGs is not complete, with an estimated $\sim$5,000 RSGs predicted to exist within our Galaxy [54]. nEXO’s sensitivity extends roughly to the Galactic Center of the Milky Way. We calculate there will nominally be one or more electron-neutrino interactions in nEXO for CCSNe out to $\sim$5-8 kpc, depending on the supernova neutrino flux model and parameters of the CCSN. Within that distance of Earth there are several hundred candidate RSGs that may produce neutrino-xenon interactions in nEXO. A 300-metric-ton detector of similar enrichment would be able to observe events out to significantly larger distances, with $\sim$15-41 events predicted at a distance of 10 kpc depending on which supernova electron-neutrino flux model is used.
@@
Figure 2: The number of supernova electron-neutrino interactions ($\nu_{e}-{}^{134}$eCC, $\nu_{e}-{}^{136}$eCC and $\nu_{e}-e^{-}$) occurring in nEXO and a 300-metric-ton detector of the same enrichment as a function of distance. No detector efficiencies are taken into account. The red line shows the cumulative number of candidate RSGs from Ref. [53]. nEXO’s expected range extends out to approximately the Galactic Center (black dashed line), which covers the majority of the identified RSG candidates in that survey, although the expected number in our Galaxy is much larger [54]. The supernova flux models correspond to those from Refs. [50, 51], and a pinched-thermal spectrum as described in Sec. V.1.
@@ III.5 Exclusive cross section predictions
When a neutrino interacts with a nucleus, various particles (predominantly neutrons and $\gamma$ rays) can be emitted as a result of the deexcitation of the product nucleus. MARLEY generates predictions not only for inclusive cross sections, but also for the various exclusive cross sections and particle distributions resulting from $\nu_{e}$-Xe interactions. Exclusive cross sections leading to bound cesium states (no nucleon emission), along with those resulting in neutron emission, are shown in Table 2. The dominant exclusive deexcitation channel is expected to be neutron emission, although recent measurements of CC neutrino interactions on 127I and Pb with $\sim$30 MeV neutrinos have suggested this channel may be suppressed for heavy nuclei [28, 29].
@@
Channel GKVM [50] Livermore [51] Pinched thermal
@@
Channel GKVM [50] Livermore [51] Pinched thermal
@@
Channel GKVM [50] Livermore [51] Pinched thermal
@@
Table 2: Calculated exclusive electron-neutrino cross sections on xenon from MARLEY, using the same assumptions as described in the caption of Table 1. For neutrino-electron elastic scattering, this cross section is given per xenon atom.
@@ III.6 Predicted particles and visible energy
Using MARLEY, the fraction of incident neutrino energy transferred to various output channels is visualized in Fig. 3. This plot incorporates $\nu_{e}-{}^{134}$eCC, $\nu_{e}-{}^{136}$eCC, and $\nu_{e}-e^{-}$ interactions, with appropriate cross section weighting and isotopic abundances (90% 136Xe, 10% 134Xe). Individual distributions for these three processes can be found in Appendix B.
@@
In the interactions modeled, some energy will be lost to nonvisible channels (thresholds for CC interactions, neutron-binding energy, and the scattered neutrino from elastic neutrino-electron scattering interactions). Additionally, emitted neutrons and the recoiling xenon nucleus also produce scintillation light and ionization charge, although these signals are quenched so their contribution is small [55]. To approximate the energy that would be visible in a detector from electron-neutrino interactions (neglecting detector thresholds, geometric size, and efficiencies), the distribution of visible (scintillation) energy is plotted along with the incident neutrino energy in Fig. 4, using an incident electron-neutrino spectrum characterized by the GKVM supernova flux model [50]. Here, visible energy ($E_{\mbox{vis}}$) is defined as the energy of resulting $\gamma$ rays and electrons,
@@
Figure 4: Comparison of the energy of interacting electron neutrinos and the corresponding visible energy spectrum using the GKVM flux model [50]. Neutrino-electron elastic scattering from $\bar{\nu}_{e}$ and $\nu_{x}$ are not included in this plot.
@@ IV Simulation and event reconstruction
Electron-neutrino interaction events were simulated using Geant4 v10.7.2 [56, 57, 58] along with a modified version of the Noble Element Simulation Technique 2.0.1 [59, 60]. For more details on the simulation framework, refer to Ref. [19].
@@
Second, the use of charge as an energy estimator would require a stricter fiducial-volume cut to select only charge events fully contained in the TPC drift region. As multiple MeV-scale $\gamma$’s can be emitted by a CC event as part of the cesium deexcitation process, CC events are spatially larger than the events typically of interest to nEXO. Using light as the energy estimator allows for the inclusion of events that extend into the xenon space outside the drift region (referred to as the xenon skin).
While we use light as our energy estimator, we reconstruct the position and magnitude of charge depositions to use a simulated map of the geometric dependence of light collection efficiency across the detector (see Ref. [19]) to correct for the position dependence of light collection. The impact of both grouping a light signal with an incorrect ionization signal or using partially contained charge depositions to apply this correction is a small increase in the energy resolution of our reconstructed light signals using this algorithm. In the future, an improved charge timing and energy reconstruction can be developed for nEXO supernova events, though for this study we choose to start with the above simplifications.
@@ IV.1 Photon reconstruction
For scintillation light reconstruction, the number of collected OPs was calculated based on the true interaction position, the number of produced photons, a map of the simulated geometric dependence of the photon collection efficiency, the photon detection efficiency of the SiPMs, and the correlated avalanche probability (based on values measured for prototype nEXO SiPMs [19, 61]). As in Ref. [19], dark counts and electronic noise were neglected due to their small expected contribution [62, 63]. The number of OPs was calculated in a moving 300-ns window (approximately 10 times the mean decay time of the long component of xenon scintillation [64]), and if the sum in that window exceeded a threshold (in terms of collected OPs, corresponding to $\sim$100efficiency for 500 keV events within the drift region), the integrated number of OPs within that window was recorded.
SiPM saturation was not included in the reconstruction algorithm, given the large spatial distribution of the energy depositions in CC events and isotropy of scintillation light production. The nEXO photon readout electronics are set to clip SiPM channels with more than 100 collected photoelectrons (PEs). A subset of $\nu_{e}$-Xe events were simulated using CHROMA [65, 66] to study collected photon distributions in SiPM channels. While 55.6% of reconstructed supernova neutrino events had at least one SiPM channel collecting more than 100 PEs, for events where clipping was present the mean fraction of clipped SiPM channels was 1.67%. For these events, nonclipped quantities, such as the tail integral [67] or nonsaturated SiPM channels, can be used to estimate event energy at the cost of an increase in energy resolution. Additional work is needed to quantify the impact of these mitigating approaches on energy resolution, and we have not included this effect in our study.
@@ IV.2 Charge reconstruction
The number of simulated drifted ionization electrons collected by charge tiles is calculated, including effects of electron lifetime attenuation as well as longitudinal and transverse dispersion. Similar to the light reconstruction algorithm, a rolling window of 120 $\mu$s is used to sum the number of detected electrons in each individual charge channel; if these exceed the threshold of 500 electrons, approximately 2.5 times the rms channel noise, the integrated charge in that channel is recorded.
Noise is added to the integrated charge on each channel based on a distribution derived using an optimum filter [68, 69] for reconstruction based on an internal nEXO collaboration study. This was found to be in good agreement with the data, and produces an energy resolution near the $0\nu\beta\beta$ $Q$ value consistent with that of Ref. [19]. A small number of events produce charge on channels that exceed the 120,000 electrons used in that study, so a $1/A$ dependency was fit to the channel-integrated charge noise for lower-energy charge deposits and used to extrapolate noise for higher single-channel charge deposits, where $A$ is the integral of the pulse.
The potential of having multiple clusters of charge deposition (e.g. prompt $\gamma$ ray and electron depositions plus delayed neutron captures) separated in both space and time drives the need for a clustering algorithm to group charge depositions correlated with a photon signal. The input to the clustering algorithm were charge-weighted points, $Q_{x_{i},y_{j}}$, located at the intersections of x and y strips on each charge tile,
@@
where $Q_{x_{i}}$ is the charge on channel $x_{i}$, $Q_{y_{j}}$ is the charge on channel $y_{j}$, and $Q_{x_{tile}}$ and $Q_{y_{tile}}$ are the sum of the charge on the x and y channels of the tile respectively. Then, the DBSCAN algorithm [70, 71] implementation of scikit-learn [72] is used to group individual charge points to form charge clusters. The DBSCAN algorithm has two free parameters: eps, which corresponds to the maximum distance between two points in the cluster for them to be considered within the same neighborhood, and min_samples, which corresponds to the minimum number of samples in a neighborhood to consider a point to be a core point. min_samples was left at its default value of 5, and eps was set to a value of 1,000 mm for this analysis. This is a conservative value chosen empirically by comparing the number of reconstructed charge clusters to the number of reconstructed light clusters. The purpose of the clustering algorithm is not to identify individual particles produced by electron-neutrino interactions, but instead to group interactions as part of either the prompt signal or the delayed neutron capture. The optimized large value likely originates from the broad spatial distribution of the emitted $\gamma$ rays and electron from these events, as well as scatters of the emitted neutron that can occur far from the electron-neutrino interaction location.
Charge-channel saturation was not incorporated into the reconstruction algorithm. Assuming a 12-bit digitizer is used for the charge readout, a baseline located at one-third of the digitizer range, and a gain setting corresponding to $\sim$43 $e$/analog-to-digital converter counts, it was determined that $\sim$0.04of events have at least one charge channel that would be saturated. The CC events simulated are high in energy, but typically have broad spatial distributions, owing to the production of a number of deexcitation $\gamma$’s, so the saturation of an individual charge channel occurs infrequently. Additionally, as with light signals, it may be possible to use unsaturated portions of charge signals when saturation is present. These assumptions are motivated by simulations of the optimal settings for the CRYO ASIC charge digitization chip planned for use in nEXO [73, 74].